Consider a slide shaped like a right triangle. The base of the slide is 45 feet away from the ladder. The top of the slide is 20 feet high. a) What is the slope of the slide. b) write an equation for the line that the slide would form. c) Lets say you were standing at the bottom of the slide and walked 27 feet closer to the ladder. How high is the slide at this point? d)If the slide was 50 feet long, what would be the new slope
Answer by stanbon(75887)

Consider a slide shaped like a right triangle. The base of the slide is 45 feet away from the ladder. The top of the slide is 20 feet high.
a) What is the slope of the slide.
Ans:: slope = 20/45 = 4/9
-------------------------
b) write an equation for the line that the slide would form.
slope = 4/9 ; Let the top be at (0,20)
Equation: h(x) = (4/9)x + 20
-------------------------------------
c) Lets say you were standing at the bottom of the slide and walked 27 feet closer to the ladder. How high is the slide at this point?
It is still 20 feet. Why would it change?
-------------------------------------
d)If the slide was 50 feet long, what would be the new slope
height = 20 ; base = sqrt[50^2-20^2] = sqrt(2100) = 10sqrt(21)
Slope = 20/(10sqrt(21)) = 2sqrt(21)/21
-------------------------
Cheers,
Stan H.
--------------